Pete is 1/3 of Brian’s age. Philip is 5 years older than Brian. The sum of their ages is 40. Find the age of each.
Answer by cougar12(1)

Pete=x
Phillip=y
Brian=z
x=1/3z
y=z+5
x+y+z=40
now change all to z
(1/3z)+(z+5)+z=40

multiply each by 3 to cancel out the 1/3z
now you have
3z+15+3z=105
now subtract 15 you have
3z+3z=90
now divide all by 3
z+z=30
2z=30
z=15
x=1/3z
x=1/3(15)
x=5
y=z+5
y=15+5
y=20
x+y+z=40
5+20+15=40